OpenStudy (anonymous):

hey you guys I need but I want you to explain it to me don't just give me an answer an insurance salesperson earns a salary of $4,200 per month pulse a commission of 3% of sales. how much must the salesperson sell to have a monthly income of at least $4,500?

OpenStudy (amistre64):

4500 = 4200 + 3%C should be the setup

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

subtract the 42 and divide by 3% to determine the comminssions

OpenStudy (anonymous):

The salesperson gets 4200 as his salary, and that doesn't depend on his sales. Start with that. Then, add in a percentage of his sales. What is 3 percent of a number? Can you multiply his sales by a number to get the percentage? If so, you can add his base salary to his sales percentage to get his income.

OpenStudy (anonymous):

@amistre64 okay then what I get would be the anwer?

OpenStudy (amistre64):

what you get would then be the answer yes just remember that 3% is equal to .03 :) (4500-4200)/.03 = C

OpenStudy (anonymous):

INCOME = BASE PAY + (PERCENTAGE/100) * SALES

OpenStudy (anonymous):

okay thanks you guys

OpenStudy (anonymous):

would the answer be 14

OpenStudy (amistre64):

um, no 3% of 14 is like .... 42 cents.

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

4500-4200 = 300 300/.03 is in the tens of thousands i think

OpenStudy (anonymous):

oh oka

OpenStudy (anonymous):

okay